OpenStudy (anonymous):

Amanda is 12 years old and is 2/7 times her fathers age. how many years ago was her father 4 times as old as her? Explain

OpenStudy (anonymous):

I need like why this answer is right and explaining the thinking

OpenStudy (zzr0ck3r):

1) Amanda is 2/7 of her fathers age \[2) \ \ Amandas\ \ age = \frac{2}{7}fathers \ age\] lets think why these two things are the same 50 cents is 1/2 of 1 dollar \[.50=\frac{1}{2}1\] ok so now we just need to plug in amandas age and then solve for fathers age \[12=\frac{2}{7}f\ \ \text{where f is her father}\] does this make sense?

OpenStudy (zzr0ck3r):

brb phone...

OpenStudy (anonymous):

yeah....... but I need some thing leas complicated

OpenStudy (zzr0ck3r):

it does not get less complicated than 12 = 2/7 x

OpenStudy (anonymous):

*less

OpenStudy (jdoe0001):

she's 2/7 the age of her father today, she's 12 so we know that 12 is 2/7 of HIS age, that is f = father a = alice \(\bf a = 12\qquad a = \cfrac{2}{7}f\\ \quad \\ 12 = \cfrac{2}{7}f\) solve for "f" to get HIS age today

OpenStudy (anonymous):

no the explanation is complicated

OpenStudy (zzr0ck3r):

lol in what way is this any different than what I said?

OpenStudy (jdoe0001):

heheh it isn't =)

OpenStudy (zzr0ck3r):

seems pointless and more confusing then for the user?

OpenStudy (anonymous):

im being taught 5 grades ahead!!!

OpenStudy (jdoe0001):

hmmm... ahem.. I see

OpenStudy (anonymous):

im in1st grade

OpenStudy (anonymous):

im happy I can do quadratics though

OpenStudy (jdoe0001):

hmm.... so.... seems to me this exercise may not apply to you then, since it involves fractions

OpenStudy (anonymous):

i know fractions!!!!!!!!!!!!!!

OpenStudy (anonymous):

ohhhh ok i get the explanation now!

OpenStudy (anonymous):

move on

OpenStudy (jdoe0001):

\(\bf a = 12\qquad a = \cfrac{2}{7}f\\ \quad \\ a = a\implies 12 = \cfrac{2}{7}f\)

OpenStudy (anonymous):

yes i saw that

OpenStudy (jdoe0001):

well, anyhow, what would be his age today then?

OpenStudy (anonymous):

42 years old

OpenStudy (anonymous):

gah! lag!!

OpenStudy (jdoe0001):

so, he's 42 today, she's 12 how many years ago was she 1/4 his age, or he was 4 times her age? well, she's 12, we can only go back 12 years, otherwise she wasn't born yet =) so, when she was 11, he was 41 is 11 * 4 = 14? no so, when she was 10, he was 40 what do you think?

OpenStudy (anonymous):

yesss!!!

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

11 * 4 = 41 doh. anyhow.. just a typo,, but yes when she's 10, he's 4 times 10

OpenStudy (anonymous):

il get problems after this I GOT 100 PROBLEMS

OpenStudy (anonymous):

GAWD LUV THIS WEBSITE